solve using long division, then check.
(x^3+2x^2+3x-6)/(x-1)=?

Answer by lwsshak3(11628)

solve using long division, then check.
(x^3+2x^2+3x-6)/(x-1)=?
**
..............x^2......3x.......6.......
...x-1|....x^3....2x^2....3x...-6
........................3x^2....3x......
........................3x^2..-3x......
....................................6x..-6..
....................................6x..-6..
..........................................0..
Check:
quotient*divisor+remainder=dividend
(x^2+3x+6)(x-1)+0
=x^3+3x^2+6x-x^2-3x-6+0
=x^3+2x^2+3x-6